OpenStudy (jskinny0831):

The value V(t) of a $42,000 car after t years that depreciates at 15% per year is given by the formula V(t) = V0(b)t. What is the value of V0 and b, and what is the car's value after 5 years (rounded to the nearest cent)? Hint: Use the equation b = 1 − r to determine depreciation. A. V0 = $42,000, b = 0.85, and the value after 5 years is $35,700.00 B. V0 = $42,000, b = 0.15, and the value after 5 years is $3.19 C. V0 = $42,000, b = 0.85, and the value after 5 years is $18,635.62 D. V0 = $42,000, b = 1.15, and the value after 5 years is $84,477.00

OpenStudy (jskinny0831):

@soprano.h.d0816

OpenStudy (jskinny0831):

@RhondaSommer

OpenStudy (jskinny0831):

please help me

OpenStudy (soprano.h.d0816):

Which do you think it is?

OpenStudy (jskinny0831):

i was gonna put D but idk

OpenStudy (soprano.h.d0816):

That sounds about right now that I think about it

OpenStudy (jskinny0831):

really?

OpenStudy (soprano.h.d0816):

Yeah. At first I thought it was C.

OpenStudy (jskinny0831):

why'd you think it was C

OpenStudy (jskinny0831):

maybe I'm wrong

OpenStudy (jskinny0831):

@DanJS do you think I'm right?

OpenStudy (soprano.h.d0816):

I've seen this problem before and I thought the answer was C, I just couldn't remember how I did it first.

OpenStudy (danjs):

tha tis a general exponential function y(t) = A*b^t the value at a time t, is the original value to start 'A', times the rate of change for each time

OpenStudy (soprano.h.d0816):

So I was right? lol

OpenStudy (jskinny0831):

lol so the answer is D?

OpenStudy (danjs):

so for each year, it decreases by 15% each year, the value will become 100-15 or 85% of what it was, each increase in t year, the original value will be multiplied by 0.85

OpenStudy (danjs):

Value at year t V(t) = 42000*(0.85)^t

OpenStudy (soprano.h.d0816):

So, C

OpenStudy (jskinny0831):

A. V0 = $42,000, b = 0.85, and the value after 5 years is $35,700.00 B. V0 = $42,000, b = 0.15, and the value after 5 years is $3.19 C. V0 = $42,000, b = 0.85, and the value after 5 years is $18,635.62 D. V0 = $42,000, b = 1.15, and the value after 5 years is $84,477.00 which one then?

OpenStudy (jskinny0831):

its either A or C

OpenStudy (danjs):

yeah at year t=6, looks good

OpenStudy (danjs):

V(6) = 45000*(0.85)^6

OpenStudy (soprano.h.d0816):

C

OpenStudy (jskinny0831):

where'd you get 6 from?